Write the slope-intercept equation for the line containing the given pair of points. (−3,4) and (3,1)
Answer by Maths68(1474)

Points (−3,4) and (3,1)
Use point-slope Equation
m=(y2-y1)/(x2-x1)
m=(1-4)/(3-(-3))
m=(-3)/(3+3)
m=-3/6
m=-1/2
Equation of the line Slope-intercept form
y=mx+b
Put the value of any one point and slope we have found in above equation
m=-1/2, Point (-3,4)
4=-1/2(-3)+b
4=3/2+b
4-3/2=b
(8-3)/2=b
5/2=b
Now
m=-1/2, Point (3,1)
1=-1/2(3)+b
1=-3/2+b
1=-3/2+b
1+3/2=b
(2+3)/2=b
5/2=b
So b=5/2
Now again put the value of b and m in Equation of the Slope-intercept Form
y=mx+b
y=(-1/2)x+5/2
Above equation is our required Equation of the Straight line Slope-inercept form